OpenStudy (anonymous):

Which of these ordered pairs is a solution to the inequality 2x + y > – 4? (–3, 0) (–1, –1) (4, –12) (0, –5)

OpenStudy (anonymous):

the answer would be (-1,-1), because -1*2=-2 -2+-1=-3 which is more than -4. hope it helps :)

OpenStudy (anonymous):

thx